Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this.
8y + 2x = 48
2x = 4y – 12

Answer by lefty4ever26(59)

You need to solve one of the problems for one of the variables.
I chose to solve the 2nd equation for x.
2x=4y-12 Divide both sides by 2.
x=2y-6 Now you have the value of x. Plug (2y-6) into your x value in the first equation and solve for y.
8y+2x=48
8y+2(2y-6)=48
8y+4y-12=48
12y-12=48 Add 12 to both sides
12y=60 Divide both sides by 12
y=5 So we have the y value. Plug it back into one of the equations to solve for x. I chose the second equation
2x=4y-12
2x=4(5)-12 Solve for x
2x=20-12
2x=8
x=4
So the solution to the system of equations is (4,5).
Hope this helps!